# Christian Shelton, Manmohan Chandraker‡,⋆, Abhishek Aich

### v1

@@ 4 Experiments
We qualitatively analyze all methods in Figure˜3 and observe the following. One, $\bm{\dot{\imath}}$Finder demonstrates superior perceptual grounding—for instance, in the first case, unlike VideoLLaMA2 and VideoChat2, $\bm{\dot{\imath}}$Finder correctly identifies the pedestrian as an immediate hazard, aligning with the ground truth. Two, $\bm{\dot{\imath}}$Finder shows fine-grained causal reasoning: in the second row, $\bm{\dot{\imath}}$Finder specifies that the ego vehicle reduced its distance from 24.75m to 6.33m, clearly linking this to the collision—something missing in the generic response by VideoLLaMA2. Three, $\bm{\dot{\imath}}$Finder exhibits higher temporal precision, as seen in the final example where $\bm{\dot{\imath}}$Finder accurately detects the collision at frame 552 (18.4s), closely matching the GT (20.167s), while others offer vague or less aligned time windows. These examples underscore $\bm{\dot{\imath}}$Finder’s robust visual grounding and decision-making fidelity.
@@
Ablation studies.
In Table˜7, we assess the impact of each vision component, as well as the contribution of each prompt component in $P_{\text{LLM}}$ on the MM-AU dataset. Although each vision component contributes to overall performance, surprisingly, scene understanding and orientation estimation are of higher importance for semantic reasoning than even core perception modules like distance or lane estimation. For example, removing orientation estimation leads to the largest drop in accuracy—from 63.39% to 58.83%. The results show that each block aid accuracy, with key explanations being especially crucial—without them, the LLM sometimes generates invalid responses, such as mis-formatted numerical outputs.
@@
Post-hoc driving video analysis also encompasses scenarios involving adversarial or long-tail conditions, such as poor weather or nighttime environments. To evaluate the robustness of our method under such conditions, we break down the performance across different weather and lighting conditions using the MM-AU[22] benchmark in Table˜5. While generalist models (e.g., VideoLLaMA2[14]) show steep drops in performance under rain (44.83%), snow (40.25%), and night (50.23%), and driving-specialized models like DriveMM perform poorly (e.g., 20.13% on snowy, 24.14% on rainy), $\bm{\dot{\imath}}$Finder maintains consistently high accuracy—achieving the best scores across all settings.
@@
With the object detector being the earliest and most important module to capture object semantics in $\bm{\dot{\imath}}$Finder, we analyze the impact of error propagation or missed detections occurring early in the system on the overall performance in Table˜7. It can be observed that even when fewer than 1% of objects are retained (0.68% at confidence score $\tau_{\text{score}}=0.8$), the accuracy remains at 58.27%, better than the best baseline (VideoLLaMA2[14] at 52.89%). At $\tau_{\text{score}}=0.5$, retaining only 32.63% of objects still yields 61.09% accuracy, suggesting that $\bm{\dot{\imath}}$Finder’s reasoning remains stable under significant perceptual filtering. Overall, $\bm{\dot{\imath}}$Finder demonstrates graceful degradation and limited error propagation.
@@ 5 Conclusion
In this paper, we argue that grounding general-purpose LLMs using structured, interpretable scene representations offers a more accurate alternative than V-VLMs. Further, decoupling perception from reasoning enables a more reliable analysis system. Built on these principles, we introduce $\bm{\dot{\imath}}$Finder, a modular, training-free framework that advances post-hoc driving video analysis by structurally grounding general-purpose LLMs in domain-specific perception. Through explicit scene decomposition and hierarchical prompting, $\bm{\dot{\imath}}$Finder achieves superior spatial and causal reasoning—outperforming both generalist and driving specialized V-VLMs on accident reasoning benchmarks. Extensive evaluations confirm strong post-hoc video understanding performance under diverse environmental conditions. This structured grounding method highlights a promising direction for aligning LLMs with domain-specific reasoning requirements, where interpretability and reliability, constitute the primary focus.
@@
Appendix A Limitations & Broader Impacts
Limitations.
The current form of $\bm{\dot{\imath}}$Finder lacks mechanisms to incorporate or reason about ambiguous, social, or normative aspects of driving scenes (e.g., intent behind a maneuver, yielding behavior, etc.). These elements are often critical in understanding traffic interactions but are not easily captured by purely spatial-temporal features or symbolic grounding. Future works should build this capability using hybrid reasoning mechanisms that combine structured perceptual data with commonsense knowledge bases.
Broader Impacts.
Appendix B Implementation choices
In Step 3, we sample the temporal points in order to reduce noise and make the estimation insensitive to small deviations. Further, we set $\tau_{a}$ and $\tau_{s}$ as 30 and standard deviation of all speeds $\{s_{t}\}_{t=0}^{T}$. For motion estimation, we set $g$ as 2. In Step 4, since we use Owl-V2, we set the 2D classes as [‘motorcycle’, ‘police car’, ‘ambulance’, ‘bicycle’, ‘traffic light’, ‘stop sign’, ‘road sign’, ‘construction worker’, ‘police officer’, ‘ambulance’, ‘fire truck’, ‘construction vehicle’, ‘traffic cone’, ‘person’, ‘car’, ‘wheelchair’, ‘bus’, ‘truck’] with confidence threshold as 0.25. In Step 5, we only estimate lane locations for vehicles and person categories. In Step 8, we use the default classes by CenterTrack [73] for NuScenes dataset [78]. All the rest of the parameters are set as default model choices. All the prompts are provided in Section F. Note that for peer V-VLM, we use the default prompt provided by respective authors. All experiments were conducted on a single NVIDIA A6000 GPU with 48 GB of memory. Table˜R1 reports the average wall-clock time required to execute each module in our pipeline per video on the LingoQA dataset [76]. These timings reflect end-to-end processing, including loading, inference, and output serialization (for unoptimized python code).
@@ Appendix B Implementation choices
Appendix C Evaluation Metrics & Dataset Details
Evaluation Metrics.
The performance on $M$-AU, SUTD, and Nexar datasets are measured using accuracy based on correct or incorrect predictions. For LingoQA, we use Lingo-Judge [76], BLEU [79], METEOR [80], and CIDEr [81]. None of the $\bm{\dot{\imath}}$Finder components are pre-trained on any of these datasets.
Dataset Details.
MMAU contains a test set of 1,953 ego-view accident videos, each associated with a fixed question: "What is the cause of the accident?" along with five multiple-choice answer options. The SUTD-TrafficQA dataset includes a test set of 4,111 real-world driving videos, paired with 6,075 multiple-choice questions designed to assess different aspects of scene understanding. The questions are categorized into six reasoning types: Basic Understanding, which involves direct perception of scene elements; Event Forecasting, which requires predicting future events; Reverse Reasoning, which focuses on deducing past events from the current scene; Counterfactual Inference, which evaluates hypothetical scenarios; Introspection, which involves providing preventive advice; and Attribution, which involve causal reasoning and responsibility assessment in driving scenarios. Performance on both datasets is measured using accuracy. For open-ended VQA, we evaluate on LingoQA [76], which consists of 100 videos with a total of 500 questions in the evaluation set. Unlike MM-AU [22] and SUTD-TrafficQA [75], which follow a multiple-choice format, LingoQA [76] requires free-form natural language responses. For accident occurrence prediction, we evaluate on the Nexar dataset [77]. Since the original test set does not include ground-truth labels, we randomly sample 100 videos from the training set to construct an evaluation set, maintaining a balanced distribution of 50 accident and 50 non-accident videos, consistent with the ratio in the full training set. The list of sampled videos from the Nexar dataset’s original training set, used for accident occurrence prediction evaluation in this paper, is provided in Table˜R2. The list indicates the video names.
@@ Appendix D Error Analysis
Figure R1: Lane location estimation. Detected objects are assigned a lane by mapping the bottom midpoint of corresponding bounding box (bottom middle point of image for ego) to sections identified by the lane detection model.
Figure R2: Distance estimation. Each object’s distance is determined by averaging the depth values within its segmented region.
Appendix E Illustration of Distance and Lane Estimation
In Figure˜R2 and Figure˜R2, we demonstrate the proposed rule to estimate distance and lane location of objects in Step 5 and Step 6, respectively.
Appendix F Prompts
The system and user prompt we use for each of the tasks in the final reasoning are shown in LABEL:lst:sys_prompt, LABEL:lst:user_prompt, LABEL:lst:sys_prompt_open_vqa, LABEL:lst:user_prompt_open_vqa, and LABEL:lst:user_prompt_acc_pred.
@@ Appendix F Prompts
Appendix G Example of Extracted Data Structure
Figure˜R3 presents an example data extracted from corresponding video (LingoQA dataset) using $\bm{\dot{\imath}}$Finder.
@@ Appendix G Example of Extracted Data Structure
Appendix H Additional Qualitative Results
We provide additional qualitative comparisons in Figure˜R4 and Figure˜R5 to further illustrate the advantages of $\bm{\dot{\imath}}$Finder over baseline methods. Figure˜R4 shows two examples where $\bm{\dot{\imath}}$Finder corrects the peer V-VLM. Figure˜R5 shows two examples where $\bm{\dot{\imath}}$Finder shows better grounded responses to user’s questions compared to baselines.

### v3

@@ 4 Experiments
We qualitatively analyze all methods in Figure˜5 and observe the following. One, $\bm{\dot{\imath}}$Finder demonstrates superior perceptual grounding—for instance, in the first case, unlike VideoLLaMA2 and VideoChat2, $\bm{\dot{\imath}}$Finder correctly identifies the pedestrian as an immediate hazard, aligning with the ground truth. Two, $\bm{\dot{\imath}}$Finder shows fine-grained causal reasoning: in the second row, $\bm{\dot{\imath}}$Finder specifies that the ego vehicle reduced its distance from 24.75m to 6.33m, clearly linking this to the collision—something missing in the generic response by VideoLLaMA2. Three, $\bm{\dot{\imath}}$Finder exhibits higher temporal precision, as seen in the final example where $\bm{\dot{\imath}}$Finder accurately detects the collision at frame 552 (18.4s), closely matching the GT (20.167s), while others offer vague or less aligned time windows. These examples underscore $\bm{\dot{\imath}}$Finder’s robust visual grounding and decision-making fidelity.
@@
Table 6: Result breakdown on MMAU weather categories. $\bm{\dot{\imath}}$Finder achieves best performance across adversarial conditions, showing its adaptability to diverse environments.
Method Foggy Rainy Snowy Sunny
Generalist Models
VideoLLaMA2 (w/pt) [damonlpsg2024videollama2] 66.67 44.83 40.25 54.56
VideoChat2 [li2023videochat] 58.33 54.31 49.69 49.16
Video-LLaVA [lin2023video] 58.33 41.38 42.77 43.76
Driving-specialized Methods
DriveMM [huang2024drivemm] 33.33 24.14 20.13 24.55
$\bm{\dot{\imath}}$Finder (Ours) 75.00 65.52 57.86 63.69
Ablation studies.
In Table˜6, we assess the impact of each vision component, as well as the contribution of each prompt component in $P_{\text{LLM}}$ on the MM-AU dataset. Although each vision component contributes to overall performance, surprisingly, scene understanding and orientation estimation are of higher importance for semantic reasoning than even core perception modules like distance or lane estimation. For example, removing orientation estimation leads to the largest drop in accuracy—from 63.39% to 58.83%. The results show that each block aid accuracy, with key explanations being especially crucial—without them, the LLM sometimes generates invalid responses, such as mis-formatted numerical outputs.
Table 7: Result breakdown on MMAU light categories. $\bm{\dot{\imath}}$Finder shows consistent improvements across different lighting conditions, indicating its robustness to variations in illumination.
Method Day Night
Generalist Models
VideoLLaMA2 (w/pt)[damonlpsg2024videollama2] 53.23 50.23
VideoChat2[li2023videochat] 50.29 43.84
Video-LLaVA[lin2023video] 43.25 46.58
AD-specialized Methods
DriveMM[huang2024drivemm] 23.41 30.59
$\bm{\dot{\imath}}$Finder (Ours) 63.32 63.93
@@
Post-hoc driving video analysis also encompasses scenarios involving adversarial or long-tail conditions, such as poor weather or nighttime environments. To evaluate the robustness of our method under such conditions, we break down the performance across different weather and lighting conditions using the MM-AU[fang2024abductive] benchmark in Table˜6 and Table˜8. While generalist models (e.g., VideoLLaMA2[damonlpsg2024videollama2]) show steep drops in performance under rain (44.83%), snow (40.25%), and night (50.23%), and driving-specialized models like DriveMM perform poorly (e.g., 20.13% on snowy, 24.14% on rainy), $\bm{\dot{\imath}}$Finder maintains consistently high accuracy—achieving the best scores across all settings.
@@
With the object detector being the earliest and most important module to capture object semantics in $\bm{\dot{\imath}}$Finder, we analyze the impact of error propagation or missed detections occurring early in the system on the overall performance in Table˜8. It can be observed that even when fewer than 1% of objects are retained (0.68% at confidence score $\tau_{\text{score}}=0.8$), the accuracy remains at 58.27%, better than the best baseline (VideoLLaMA2[damonlpsg2024videollama2] at 52.89%). At $\tau_{\text{score}}=0.5$, retaining only 32.63% of objects still yields 61.09% accuracy, suggesting that $\bm{\dot{\imath}}$Finder’s reasoning remains stable under significant perceptual filtering. Overall, $\bm{\dot{\imath}}$Finder demonstrates graceful degradation and limited error propagation.
@@ 5 Conclusion
In this paper, we argue that grounding general-purpose LLMs using structured, interpretable scene representations offers a more accurate alternative than V-VLMs. Further, decoupling perception from reasoning enables a more reliable analysis system. Built on these principles, we introduce $\bm{\dot{\imath}}$Finder, a modular, training-free framework that advances post-hoc driving video analysis by structurally grounding general-purpose LLMs in domain-specific perception. Through explicit scene decomposition and hierarchical prompting, $\bm{\dot{\imath}}$Finder achieves superior spatial and causal reasoning—outperforming both generalist and driving specialized V-VLMs on accident reasoning benchmarks. Extensive evaluations confirm strong post-hoc video understanding performance under diverse environmental conditions. This structured grounding method highlights a promising direction for aligning LLMs with domain-specific reasoning requirements, where interpretability and reliability constitute the primary focus.
6 Acknowledgments and Disclosure of Funding
This work was supported by NSF grant CNS-2312395 and NEC Laboratories America, Inc.
@@ 6 Acknowledgments and Disclosure of Funding
Appendix A Limitations & Broader Impacts
Limitations.
The current form of $\bm{\dot{\imath}}$Finder lacks mechanisms to incorporate or reason about ambiguous, social, or normative aspects of driving scenes (e.g., intent behind a maneuver, yielding behavior, etc.). These elements are often critical in understanding traffic interactions but are not easily captured by purely spatial-temporal features or symbolic grounding. Future work should build this capability using hybrid reasoning mechanisms that combine structured perceptual data with commonsense knowledge bases.
Broader Impacts.
Appendix B Implementation choices
In Step 3, we sample the temporal points in order to reduce noise and make the estimation insensitive to small deviations. Further, we set $\tau_{a}$ and $\tau_{s}$ as 30 and standard deviation of all speeds $\{s_{t}\}_{t=0}^{T}$. For motion estimation, we set $g$ as 2. In Step 4, since we use Owl-V2, we set the 2D classes as [‘motorcycle’, ‘police car’, ‘ambulance’, ‘bicycle’, ‘traffic light’, ‘stop sign’, ‘road sign’, ‘construction worker’, ‘police officer’, ‘ambulance’, ‘fire truck’, ‘construction vehicle’, ‘traffic cone’, ‘person’, ‘car’, ‘wheelchair’, ‘bus’, ‘truck’] with confidence threshold as 0.25. In Step 5, we only estimate lane locations for vehicles and person categories. In Step 8, we use the default classes by CenterTrack [zhou2020tracking] for NuScenes dataset [caesar2020nuscenes]. All the rest of the parameters are set as default model choices. All the prompts are provided in Section E. Note that for peer V-VLM, we use the default prompt provided by the respective authors. All experiments were conducted on a single NVIDIA A6000 GPU with 48 GB of memory. Table˜R1 reports the average wall-clock time required to execute each module in our pipeline per video on the LingoQA dataset [marcu2024lingoqa]. These timings reflect end-to-end processing, including loading, inference, and output serialization (for unoptimized python code).
Appendix C Dataset Details
@@
Dataset Details.
MMAU contains a test set of 1,953 ego-view accident videos, each associated with a fixed question: "What is the cause of the accident?" along with five multiple-choice answer options. The SUTD-TrafficQA dataset includes a test set of 4,111 real-world driving videos, paired with 6,075 multiple-choice questions designed to assess different aspects of scene understanding. The questions are categorized into six reasoning types: Basic Understanding, which involves direct perception of scene elements; Event Forecasting, which requires predicting future events; Reverse Reasoning, which focuses on deducing past events from the current scene; Counterfactual Inference, which evaluates hypothetical scenarios; Introspection, which involves providing preventive advice; and Attribution, which involve causal reasoning and responsibility assessment in driving scenarios. Performance on both datasets is measured using accuracy. For open-ended VQA, we evaluate on LingoQA [marcu2024lingoqa], which consists of 100 videos with a total of 500 questions in the evaluation set. Unlike MM-AU [fang2024abductive] and SUTD-TrafficQA [Xu_2021_CVPR], which follow a multiple-choice format, LingoQA [marcu2024lingoqa] requires free-form natural language responses. For accident occurrence prediction, we evaluate on the Nexar dataset [nexar_dataset]. Since the original test set does not include ground-truth labels, we randomly sample 100 videos from the training set to construct an evaluation set, maintaining a balanced distribution of 50 accident and 50 non-accident videos, consistent with the ratio in the full training set. The list of sampled videos from the Nexar dataset’s original training set, used for accident occurrence prediction evaluation in this paper, is provided in Table˜R2. The list indicates the video names.
Appendix D Error Analysis
Although $\bm{\dot{\imath}}$Finder achieves strong results, it struggles in visually ambiguous collisions. For example, from the Nexar dataset, video “01031”, a car cuts in front of the ego vehicle, but without clear cues such as steering correction or vibration, even humans cannot confirm contact; this reveals $\bm{\dot{\imath}}$Finder ’s reliance on observable patterns rather than imperceptible dynamics. In the video “00970”, the ego car brakes sharply behind another vehicle, leaving only a small gap. With no visible impact signs like deformation or debris, it is unclear whether this was a collision or a near-miss. Both cases show that subtle physical contact often leaves no visual trace, limiting vision-only approaches. Future work should integrate other modalities (e.g., audio, IMU) and model uncertainty so the system can express lower confidence instead of making forced binary predictions.
Appendix E Prompts
The prompts we use for the VLM are shown in LABEL:lst:prompt_i, LABEL:{lst:prompt_v} and LABEL:lst:prompt_d. The system and user prompts we use for each of the tasks in the final reasoning are shown in LABEL:lst:sys_prompt, LABEL:lst:user_prompt, LABEL:lst:sys_prompt_open_vqa, LABEL:lst:user_prompt_open_vqa, and LABEL:lst:user_prompt_acc_pred.
@@ Appendix E Prompts
Appendix F Example of Extracted Data Structure
Figure˜R1 presents an example of data extracted from the corresponding video (LingoQA dataset) using $\bm{\dot{\imath}}$Finder.
@@ Appendix F Example of Extracted Data Structure
Appendix G Additional Qualitative Results
We provide additional qualitative comparisons in Figure˜R2 and Figure˜R3 to further illustrate the advantages of $\bm{\dot{\imath}}$Finder over baseline methods. Figure˜R2 shows two examples where $\bm{\dot{\imath}}$Finder corrects the peer V-VLM. Figure˜R3 shows two examples where $\bm{\dot{\imath}}$Finder shows better grounded responses to users’ questions compared to baselines.